# A recursive approach for the determination of the nice sections of width three which have a 4-crown stack as retract Version 8-3

### v2

@@ 4.1 Definition and isomorphism types
For each integer $n\geq 2$ and each index set $K\subseteq\{0,\ldots,n-1\}$ with $0,n-1\in K$, we can construct a poset $P\in\mathfrak{N}_{2}$ with $h_{P}=n$ and $P(k,k+1)$ being a 6-crown iff $k\in K$: We start with the union $Q$ of three pairwise disjoint chains of length $n$, extend $Q(k,k+1)$ to an arbitrary 6-crown for every $k\in K$, and complete the partial order relation by adding $Q(k)\times Q(\ell)$ for all $k,\ell\in[0,n]$ with $k\leq\ell-2$. The resulting section is an element of $\mathfrak{C}_{2}$ and belongs thus to $\mathfrak{N}_{2}$ (Proposition 4.1).
@@
We denote the classes of segments, lower segments, and upper segments of nice sections with horizon 2 by $\mathfrak{N}^{Seg}_{2},\mathfrak{N}^{LS}_{2}$, and $\mathfrak{N}^{US}_{2}$, respectively. Theorem 4.2 extends to these classes: Two segments of equal height $n$ are isomorphic iff both are 3-antichains (the case $n=0$) or all of their consecutive level-pairs are isomorphic (the case $n\geq 1$), and also the extension of a level permutation to an automorphism is always possible.
@@ 4.2 Properties of retracts
The first proposition follows from the definition of the horizon and (1). For the second one, observe that the assumption implies $a\leq r(y)$ for all $a\in R(\ell)$ and all $y\in P(\rho+1)$, hence $R(\ell)<r(y)$.
@@ 4.3 Retractive up- and down-splits
Let $P\in\mathfrak{N}_{2}$ and let $r:P\rightarrow R$ be a retraction onto a 4-crown stack $R$. Due to the horizon of $P$, every level set of $R$ is either contained in a single level set of $P$ or consists of points belonging to consecutive level sets of $P$. Now Lemma 3.4 and Lemma 3.6 tell us that every level-set of $R$ gives raise to matching retractive split of $P$. On the other hand, for a given retractive split of $P$ it is easy to check if there exists a matching retraction of $P$ onto a 4-crown stack:
@@
Let $r:P\rightarrow R$ be a retraction onto a 4-crown stack $R$ matching a retractive up-split $(k,U,s,t)$. Due to $r[P(k\rightarrow h_{P})]=T$, the second part of Condition (4) is necessary for $r[U]\subseteq T$. The addendum follows because of $P(k-2)<_{P}P(k)\supset T(0)$.
@@
Figure 6: Illustrations for the proof of Corollary 4.6. The three levels show $P(k-1\rightarrow k+1)$ for choices $b\in P(k-1)$, $w\in P(k+1)$ (which are not mandatory). The point $z$ is not mapped to $a$ by the retraction $s$ of $P(0\rightarrow k-1)$.
@@
The 4-crown stack $S\oplus T$ is a retract of $P$ if $P(k-1\rightarrow k+1)$ is not a 6-crown stack.
@@
Let $S(h_{S})=\{a,b\}$ and $T(0)=\{v,w\}$. The assumptions imply $a\in P(k-1)$ and $v\in P(k+1)$. Assume $P(k,k+1)\simeq 3C$ (the other case is dual). According to Lemma 2.1, we can assume $S(h_{S})\subset P(k-1)$ without loss of generality. Let $P(k-1)=\{a,b,z\}$.
Let $U\subset P(k)$ be the set of upper covers of $a$ in $P(k)$. Applying Theorem 4.2 on $P(0\rightarrow k)$, we see that without loss of generality we can assume that none of the points in $U$ is below $v$ (Figure 6).
@@ 5 Application
With the recursive approach described in Section 3.3, we will now identify all posets $P\in\mathfrak{N}^{LS}_{2}$ with height up to six which have a 4-crown stack as retract. For some of them, e.g. the 6-crown stacks, we already know the result. We include them in the investigation because our purpose is to test our approach with as many nice sections as possible.
@@
    $k=0$ and $P(1,2)$ being a 6-crown implies that the poset $P(3\rightarrow h_{P})$ has a 2-antichain or a 4-crown stack as retract.
@@
    $k=h_{P}-2$ and $P(h_{P}-2,h_{P}-1)$ being a 6-crown implies that the poset $P(0\rightarrow h_{P}-3)$ has a 2-antichain or a 4-crown stack as retract.
@@
Figure 9: The posets $1111$, $1101$, $1011$, and $1001$. As already mentioned at the end of the introduction, the poset $1011$ is also treated in [9, p. 99] and mentioned in [3, p. 136] as a very complicated case. For the poset $1001$, a retraction onto a 4-crown tower is indicated resulting from a retractive down-split $(2,D,s,t)$ based on the retraction $t$ of $P(0\rightarrow 2)=10$ shown in Figure 7a.
@@
  • $P=11111$: Due to the Criteria 5.1.2 and 5.1.4, the upper segments $P(1\rightarrow 5)$ and $P(4,5)$ cannot be work as $s$-bases. The self-dual poset has thus no 4-crown stack as retract.
@@
  • $P=10101,10001$: Let $s:P(0\rightarrow 2)\rightarrow S$ and $t:P(3\rightarrow 5)\rightarrow T$ retractions onto 4-crowns as shown in Figure 7a and its dual. Without loss of generality (Theorem 4.2) we assume that the single point in $P(2)\cap S$ is below the single point in $P(3)\cap T$. Now $(3,\emptyset,s,t)$ is a retractive down-split fulfilling (3).
@@
Firstly, let $P\in\mathfrak{N}_{2}$ be any nice section of height six and $R$ a retract of $P$ being a 4-crown stack. Every level set of $R$ has to be contained in a single level set or in two consecutive level sets of $P$, and we conclude that at least one of the posets $P(0\rightarrow 3)\cap R$ and $P(3\rightarrow 6)\cap R$ is a 4-crown stack.
Now let $P$ be any of the four posets and $r$ a retraction onto a 4-crown stack $R$. With $S:=P(3\rightarrow 6)\cap R$, we show in a first step that $S$ cannot be a 4-crown stack.
$P(0\rightarrow 3)=111,110$: $P(3\rightarrow 6)$ does not have a 4-crown stack as retract. $S$ being a 4-crown stack requires thus a point $x\in P(3)$ with $v:=r(x)\in P(0\rightarrow 2)=11$. Let $\{v,w\}$ be the level set of $R$ containing $v$. $x\in P(3)$ yields $r[P(0,1)]\leq v$, hence $w\notin P(0,1)$. But then, $r(z)<\{v,w\}$ holds for at least five of the points $z\in P(0,1)$, and $P(0,1)$ cannot be mapped onto a 2-antichain.
$P=101011$: $S$ being a 4-crown stack requires due to Lemma 5.2 two points $x,y\in P(3)$ to be mapped into $P(0\rightarrow 2)=10$. The points $r(x)$ and $r(y)$ belong thus to $R(0,1)$. Three cases are possible, each leads to a contradiction:
@@
  • $r(x)<r(y)$: Then $r(x)\in R(0)$, and $P(0)<x$ yields that $r(x)$ is the only minimal point of $R$.
  • $v:=r(x)=r(y)$: Let $\{v,w\}$ be the level set of $R$ containing $v$. Due to $r[P(0\rightarrow 2)]\leq v$, the point $w$ must belong to $P(3\rightarrow h_{P})$, and due to $\lambda(v)\leq 2$, we must have $w\in P(3)$ and $v\in P(2)$. But then $w\in S\not\ni v$.
$S$ is thus not a 4-crown stack. This requires a level set $R(\ell)=\{a,b\}$ with $a\in P(2)$ and $b\in P(3)$. Clearly, $\ell=1$. Due to $P(1)<b$, no point of $P(1)$ can be mapped to $a$. And if a point of $P(1)$ is mapped to $b$, (2) yields $R(2\rightarrow h_{R})$ being a retract of $P(4\rightarrow 6)=11$ which is impossible. But if no point of $P(1)$ can be mapped to $R(1)$, then $r[P(0,1)]=R(0)$ which is impossible, too.

### v3

@@ 4.1 Definition and isomorphism types
For each integer $n\geq 2$ and each index set $K\subseteq\{0,\ldots,n-1\}$ with $0,n-1\in K$, we can construct a poset $P\in\mathfrak{N}_{2}$ with $h_{P}=n$ and $P(k,k+1)$ being a 6-crown iff $k\in K$. We start with the union $Q$ of three pairwise disjoint chains of length $n$, extend $Q(k,k+1)$ to an arbitrary 6-crown for every $k\in K$, and complete the partial order relation by adding $Q(k)\times Q(\ell)$ for all $k,\ell\in[0,n]$ with $k\leq\ell-2$. The resulting section is an element of $\mathfrak{C}_{2}$ and belongs thus to $\mathfrak{N}_{2}$ (Proposition 4.1).
@@
We denote the classes of segments, lower segments, and upper segments of nice sections with horizon 2 by $\mathfrak{N}^{Seg}_{2},\mathfrak{N}^{LS}_{2}$, and $\mathfrak{N}^{US}_{2}$, respectively. Theorem 4.2 extends to these classes: Two segments of equal height $n$ are isomorphic iff both are 3-antichains (the case $n=0$) or all of their consecutive level-pairs are isomorphic (the case $n\geq 1$), and also the extension of a level permutation to an automorphism is always possible. We will also use the following fact: if $t:V\rightarrow T$ is a retraction of a segment $V\in\mathfrak{N}^{Seg}_{2}$ and $\alpha$ an automorphism of $V$, then $\alpha\circ t\circ\alpha^{-1}$ is a retraction of $V$ sharing all properties of $t$.
@@ 4.2 Properties of retracts
The first proposition follows from the definition of the horizon and (1). For the second one, observe that the assumption enforces $R(\ell)<r[P(\rho+1)]$.
@@ 4.3 Retractive up- and down-splits
Nice sections with horizon two are good objects for our recursive approach because we can expect that a 4-crown stack $R$ as retract provides “many” split levels. Let $\ell\in[1,h_{R}-1]$ be a level index of $R$. Due to the horizon of $P$, the level set $R(\ell)$ is either contained in a single level set of $P$ or in two consecutive ones. Now the first statement in Lemma 3.4 and the second part of Lemma 3.5 tell us that among the three consecutive level sets $R(\ell-1),R(\ell)$, and $R(\ell+1)$ there must be at least one split level. On the other hand, for a given retractive split of $P$ it is easy to check if there exists a matching retraction of $P$ onto a 4-crown stack:
@@
Let $r:P\rightarrow R$ be a retraction onto a 4-crown stack matching a retractive up-split $(k,U,s,t)$. Due to $r[P(k\rightarrow h_{P})]=T$, the second part of (4) is necessary for $r[U]\subseteq T$. The addendum follows because of $P(k-2)<_{P}P(k\rightarrow h_{P})\supset T(0)$.
@@
Figure 6: Illustrations for the proof of Corollary 4.6. The point $z$ is not mapped to $a$ by the retraction $s$ of $P(0\rightarrow k-1)$, and the points in $P(k)\setminus U$ are mapped to $b$ by the retraction $s^{\prime}$ of $P(0\rightarrow k)\setminus U$.
@@
$P$ has a retract isomorphic to $S\oplus T$ if $P(k-1\rightarrow k+1)$ is not a 6-crown stack.
@@
Let $S(h_{S})=\{a,b\}$ and $T(0)=\{v,w\}$. The assumptions imply $a\in P(k-1)$ and $v\in P(k+1)$. Furthermore, $s[P(k-1)\setminus\{a\}]\subseteq S\setminus\{a\}$ and $t[P(k+1)\setminus\{v\}]\subseteq T\setminus\{v\}$.
Assume $P(k,k+1)\simeq 3C$ (the other case is dual) and let $U\subset P(k)$ be the set of upper covers of $a$ in $P(k)$. Applying Theorem 4.2 on $P(k+1\rightarrow h_{P})$, we see that without loss of generality we can assume that none of the points in $U$ is below $v$. Figure 6 illustrates the case $R(\ell)\subset P(k-1)$, $R(\ell+1)\subset P(k+1)$.
@@ 5 Application
With the recursive approach described in Section 3.3, we will now identify all posets $P\in\mathfrak{N}^{LS}_{2}$ with height up to six which have a 4-crown stack as retract. For some of them, e.g. the 6-crown stacks, we already know the result. We include them in our investigation because we want to test our approach with as many nice sections as possible.
@@
    $k=0$ and $P(1,2)$ being a 6-crown imply together that the poset $P(3\rightarrow h_{P})$ has a 2-antichain or a 4-crown stack as retract.
@@
    $k=h_{P}-2$ and $P(h_{P}-2,h_{P}-1)$ being a 6-crown imply together that the poset $P(0\rightarrow h_{P}-3)$ has a 2-antichain or a 4-crown stack as retract.
@@
Figure 9: The posets $1111$, $1101$, $1011$, and $1001$. As already mentioned at the end of the introduction, the poset $1011$ is also treated in [9, p. 99] and mentioned in [3, p. 136] as a very complicated case. For the poset $1001$, a retraction onto a 4-crown tower is indicated which results from a retractive down-split $(2,D,s,t)$ based on the retraction $t$ of $P(0\rightarrow 2)=10$ shown in Figure 7a.
@@
  • $P=11111$: Due to the Criteria 5.1.2 and 5.1.4, the upper segments $P(1\rightarrow 5)$ and $P(4,5)$ cannot work as $s$-bases. The self-dual poset has thus no 4-crown stack as retract.
@@
  • $P=10101,10001$: Let $s:P(0\rightarrow 2)\rightarrow S$ and $t:P(3\rightarrow 5)\rightarrow T$ be retractions onto 4-crowns as shown in Figure 7a and its dual. Without loss of generality (Theorem 4.2) we assume that the single point in $P(2)\cap S$ is below the single point in $P(3)\cap T$. Now $(3,\emptyset,s,t)$ is a retractive down-split fulfilling (3).
@@
Firstly, let $P\in\mathfrak{N}_{2}$ be any nice section of height six and $R$ a retract of $P$ which is a 4-crown stack. Every level set of $R$ has to be contained in a single level set or in two consecutive level sets of $P$, and we conclude that at least one of the posets $P(0\rightarrow 3)\cap R$ and $P(3\rightarrow 6)\cap R$ is a 4-crown stack.
Now let $P$ be any of the four posets and $r$ a retraction onto a 4-crown stack $R$. With $Q:=P(3\rightarrow 6)\cap R$, we show in a first step that $Q$ cannot be a 4-crown stack.
$P(0\rightarrow 3)=111,110$: $P(3\rightarrow 6)$ does not have a 4-crown stack as retract. $Q$ being a 4-crown stack requires thus a point $x\in P(3)$ with $v:=r(x)\in P(0\rightarrow 2)=11$. Let $\{v,w\}$ be the level set of $R$ containing $v$. $x\in P(3)$ yields $r[P(0,1)]\leq v$, hence $w\notin P(0,1)$. But then, $r(z)<\{v,w\}$ holds for at least five of the points $z\in P(0,1)$, and $P(0,1)$ cannot be mapped onto a 2-antichain.
$P=101011$: $Q$ being a 4-crown stack requires due to Lemma 5.2 that two points $x,y\in P(3)$ are mapped into $P(0\rightarrow 2)=10$. The points $r(x)$ and $r(y)$ belong thus to $R(0,1)$. Three cases are possible, each leads to a contradiction:
@@
  • $r(x)<r(y)$: Then $r(x)\in R(0)$, and $P(0)<x$ again yields that $r(x)$ is the only minimal point of $R$.
  • $v:=r(x)=r(y)$: Let $\{v,w\}$ be the level set of $R$ containing $v$. Due to $r[P(0\rightarrow 2)]\leq v$, the point $w$ must belong to $P(3\rightarrow h_{P})$, and due to $\lambda(v)\leq 2$, we must have $w\in P(3)$ and $v\in P(2)$. But then $w\in Q\not\ni v$.
$Q$ is thus not a 4-crown stack. This requires a level set $R(\ell)=\{a,b\}$ with $a\in P(2)$ and $b\in P(3)$. Clearly, $\ell=1$. Due to $P(1)<b$, no point of $P(1)$ can be mapped to $a$. And if a point of $P(1)$ is mapped to $b$, (2) yields $R(2\rightarrow h_{R})$ being a retract of $P(4\rightarrow 6)=11$ which is impossible. But if no point of $P(1)$ can be mapped to $R(1)$, then $r[P(0,1)]=R(0)$ which is impossible, too.